OpenStudy (anonymous):

A lady reaches her office 20 minutes late by traveling at a speed of 20 kmph and reaches I5 minutes early by traveling at 30 kmph. By how much time will she be early or late if she travels at 25 kmph.

OpenStudy (anonymous):

Please help @Hero .

OpenStudy (anonymous):

hmm do u know what fomrula u have to use?

OpenStudy (anonymous):

I don't think there's any formula. You just have to use linear equations right?

OpenStudy (anonymous):

hmm let me think abt this one

OpenStudy (anonymous):

Perhaps distance = speed * time.

OpenStudy (anonymous):

well we dont know the time we just know how late they came

OpenStudy (anonymous):

How long will you be online?

OpenStudy (anonymous):

ummm idk depends how bored i get

OpenStudy (anonymous):

Can you do me a favour, if anything strikes to you about it, answer it will you? I'll have to leave now and I'll give you a medal the next time I come online ..

OpenStudy (anonymous):

hahhah dont worry abt it. i want the answer for myself. This question is annoying me

OpenStudy (anonymous):

distance = rate *time y=20(x+(20/60)) y=30(x-(15/60) 20(x+(1/3)=30(x-(1/4)) 20x+(20/3)=30x-(15/2) 10x=(85/6) x=(17/12) y=35 35=25((17/12)+z)) z=-(1/60) which is one minute early. Like i skipped some steps If u have a problem with this or need an explanation like ask and when i come back on i will explain it to u

OpenStudy (anonymous):

Let t be the amount of time to be "on time". Solve the following for t.\[20\left(t+\frac{20}{60}\right)=30\left(t-\frac{15}{60}\right)\]\[t=\frac{17}{12} \text{ hours}\]From the LHS of the equation, one can determine the distance.\[d=20\left(\frac{17}{12}+\frac{20}{60}\right)=35\text{ km} \]Have to break off for an appointment now. Back later.

OpenStudy (anonymous):

yes bu then u gotta figure out how late or early the lady wld be if she drove 25 km /hr

OpenStudy (anonymous):

To continue. From the following expression, the lady will take 35/25 hours to travel 35km to her office @ 25 km/hr.\[t\text{ =}\frac{d}{r}=\frac{35}{25}\text{hours}\]Subtract the "on time" time, 17/12, from 35/25 the time taken @ 25 km/hr. If the difference is positive the lady was "late", took too long, otherwise, the lady was "on time" or "early".\[\frac{35}{25}-\frac{17}{12}=-\frac{1}{60}\text{hours} \]The lady was one minute early @ 25 km/hr.

OpenStudy (anonymous):

YA :DDDD